OpenStudy (anonymous):

y varies directly as the square of x. When x = 5, y = 100. Find y when x = 6. y = 105 y = 120 y = 144 y = 150

OpenStudy (anonymous):

@surdawi

OpenStudy (anonymous):

y = k*x 100 = k * 5 can you calculate k?

OpenStudy (anonymous):

@rebecca1233 wanna try it?

OpenStudy (anonymous):

500

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

100 *5 = 500

OpenStudy (anonymous):

100 = k * 5 you need to divide

OpenStudy (phi):

after you figure out k for direct variation with x (as practice) you are ready to tackle the question y varies directly as the square of x \[ y= kx^2\]

OpenStudy (phi):

It is the same idea. replace x with 5 and y with 100 in the equation \[ y= kx^2\] then "solve" for k. Can you do that?